1. The plane 4x – 3y +8z = 5 intersects the cone z? = x² + y2 in an ellipse. Find the points on this ellipse the highest andplease show steps and methods used in this question
Answer #1

- Чи -3y + 8 = S , zex? у 2. We have to optimise - f(x,y,z) = z. 8( xy,z) = (х -3y +8 z-s Methad used 4 (x, y, z) = x2+y2-22.from equ (1) - 24 M Now plage these values in equ of plane 434-28) - 3.1. +4 489) = 5 -89 - 94 +4+32 = 54 397 2 ty=54 ) 39 7+- for 8 = -1/5 from equliv? 39+(-1/2) +8=10u = M = -1 / 2 for d = 1 1 1 3 from equlin) - 39% (-) +8 = 10u =) dezza. Son (9,6